OpenStudy (jbolyefo):

I'll fan and medal Find the derivative of f(x) = negative 11 divided by x at x = 9.

OpenStudy (robtobey2):

From Mathematica:\[\frac{d}{dx}\left(-\frac{11}{x}\right)\text{/.}\, x\to 9 \]\[\frac{11}{81} \]

OpenStudy (sshayer):

\[\frac{ d }{ dx}\left( \frac{ -11 }{ x } \right)=\frac{ d }{ dx }\left( -11x ^{-1} \right)=11x ^{-2}=\frac{ 11 }{ x^2 }\] put x=9 ?